OpenStudy (anonymous):

PROB&STAT ANYONE???? "Time headway" in traffic flow is the elapsed time between the time that one car finishes passing a fixed point and the instant that the next car begins to pass that point. Let X = the time headway for two randomly chosen consecutive cars on a freeway during a period of heavy flow. The pdf of X is the following. f(x) = 0.15e−0.15(x − 0.5) x ≥ 0.5 0 otherwise (a) What is the probability that time headway is at most 6 sec?

OpenStudy (anonymous):

hey nottim, are u helping me out?

OpenStudy (zarkon):

don't you just want \[\int\limits_{.5}^{6}0.15e^{−0.15(x − 0.5)}dx\]

OpenStudy (anonymous):

i dont know how to use it

OpenStudy (zarkon):

you don't know how to integrat it ?

OpenStudy (anonymous):

nope :(

OpenStudy (zarkon):

Do a substitution...let \(u=-.15(x-.5)\)

OpenStudy (anonymous):

and for x i plug in my numbers?

OpenStudy (zarkon):

you find \(du\) then you can change the limits of integration based on the subtitution \(u\)

OpenStudy (zarkon):

\[\int\limits_{.5}^{6}0.15e^{−0.15(x − 0.5)}dx\] \[u=-.15(x-.5)\] \[du=-.15dx\] \[\frac{du}{-.15}=dx\] \[\int\limits_{.5}^{6}0.15e^{−0.15(x − 0.5)}dx=\int\limits_{0}^{-.15(6-.5)}0.15e^{u}\frac{du}{-.15}\]

OpenStudy (anonymous):

what would du be?

OpenStudy (zarkon):

\[\int\limits_{0}^{-.15(6-.5)}0.15e^{u}\frac{du}{-.15}=-\int\limits_{0}^{-.825}e^{u}du\] \[-\left.e^u\right|_{0}^{-.825}=-e^{-.825}+e^{0}\approx.561765\]

OpenStudy (anonymous):

thank youuuu so much

OpenStudy (zarkon):

I hope you understand what I did.

OpenStudy (anonymous):

how'd u get -.825?

OpenStudy (anonymous):

forget i figured it out, thanks!